OpenStudy (anonymous):

Q1- 52 Cards Deck , A- How many 2 cards poker handle are possible? Q2- Fourteen horses in a race , in how many different way can first 5 finish?

OpenStudy (anonymous):

can I answer Q2 instead (sorry, not really good at cards ^^;;) use combination: 14C5 = 14!/([14-5]!5!) = (8.718x10^10)/ [(362880)(120)] =(8.718x10^10)/43545600 =2002 (not sure if i'm right though ^^;;)

OpenStudy (anonymous):

I solve it . For A- Use combination( Order is not important) = 52 C 2 = 1326 For B . this type of question the order is important . So use Permutation not combination formula 14 P 5 = 14!/(14-5)! = 240240.